QUESTION 7

  1. Katie has typically shaped indifference curves and is currently spending all of her income on jeans and sweaters. The marginal utility of the last pair of jeans was 25 and the marginal utility of the last sweater was 40. If jeans cost $30 and sweaters cost $50, then to maximize utility, Katie should:

    a.

    consume more sweaters and fewer jeans

    b.

    consumer more jeans and fewer sweaters

    c.

    Consume all sweaters as this is a corner solution.

    d.

    keep her consumption of jeans and sweaters the same

10 points   

QUESTION 8

  1. Jane's utility function is U = T + C, where T is baskets of tomatoes and C is bags of carrots. Which of the following combinations would lie on the same indifference curve for Jane?

    a.

    (T=6 and C=2) and (T=3 and C=4)

    b.

    (T=2 and C=3) and (T=1 and C=5)

    c.

    (T=6 and C=2) and (T=3 and C=5)

    d.

    (T=2 and C=2) and (T=3 and C=3)
Answer #1

7. b. consumer more jeans and fewer sweaters.
(MU of Jeans/Cost of jeans = 25/30 = 0.83
MU of Sweaters/Cost of Sweaters = 40/50 = 0.8
As MU of Jeans/Cost of jeans > MU of Sweaters/Cost of Sweaters, so more of jeans and less sweaters should be consumed.)

8. c. (T=6 and C=2) and (T=3 and C=5)
(U(6, 2) = 6 + 2 = 8
U(3, 5) = 3 + 5 = 8
So, they lie on same IC as they give same utility.)